Please help me to figure it out, thanks!

9. Use the information in the table below to answer the following questions. Unit labor requirement Food 100 180 Home Foreign Drinks 500300 Suppose Homes labor supply is 1,000 hours, and Foreigns labor supply is 2,700 hours (a) Draw the PPF for Home and for Foreign (on separate graphs). Put production of food on the x-axis.(b) Which country has the comparative advantage in food production? Which has the comparative advantage in drink production? (c) Draw the world supply of food relative to drinks (the world relative supply curve(d) Suppose that the PFood- $2/pound, and PDrinks -$5/gallon. Draw new PPFs for Home and Foreign after trade is allowed. How can we tell that trade is welfare improving in both countries?
Answer #1

a) Home can produce 1000/5 = 200 bicycles or 1000/2 = 500 skateboards. Foreign can produce 1200/3 = 400 bicycles or 1200/3 = 400 skateboards. The PPF is shown below.

b) Relative price of bicycles in Home is the opportunity cost which is 500/200 = 2.5 skateboards and in Foreign it is 400/400 = 1 skateboards. It will also be the range of terms of trade.

c) With terms of trade for bicycles being 5/4 = 1.25 skateboards. Home produces skateboards only so it will produce 500 skateboards. Foreign produces 400 bicycles. Home exports 300 skateboards and will be imports of 300/1.25 = 240 bicycles. Hence its after trade consumption is 200 skateboards and 240 cycles. In autarky, it would have produced and 200 skateboards and 120 bicycles. Now it is able to consume 120 bicycles more.

d. Foreign imports 300 skateboards and after exporting 240 bicycles it has 160 bicycles left to consume. Hence, its consumption is 300 skateboards and 160 bicycles. Without trade, it would have produced 100 bicycles and 300 skateboards. It is also better off

Production possibilities frontier Home Production possibilities frontier Foreign 400 350 300 250 200 150 100 50 400 350 300 g

Production possibilities frontier Home Production possibilities frontier Foreign 400 350 300 250 200 150 100 50 400 350 300 g